OpenStudy (anonymous):

One bicyclist rides 4 mph more than another bicyclist. The faster rider takes 3 hours to complete a race, while the slower rider takes 4 hours. Find the speed of each rider.

OpenStudy (anonymous):

Hint: This deals with having 4 added to an unknown number. Something like an "x" *hint hint*

OpenStudy (anonymous):

d=rt

OpenStudy (anonymous):

Can you get the equation setup or do you need help on that?

OpenStudy (anonymous):

i always draw a small chart w/ all the distances, times, and speeds to help me keep track of it

OpenStudy (anonymous):

I need help with the equation

OpenStudy (anonymous):

Ok here's the equation explained: You have Rider A and Rider B Rider A travels at speed x. Rider B travels at speed x+4mph. Since they are both traveling at the same distance from equation of (Distance=Rate*Time) they are equal to each other. So you have (x)mph*4 hours=(x+4)mph*3 hours Just solve for 3 from here. :D

OpenStudy (anonymous):

I meant "solve for x from here"

OpenStudy (anonymous):

good explanation Firedude

OpenStudy (anonymous):

Thanks :D, I'm just confused why the question is in the English system. Even here in the US we use the metric system in science.

OpenStudy (anonymous):

I'm a little confused with how you solve.. do you divide both sides by x?

OpenStudy (anonymous):

probably just for ease of the student. We see mph everyday.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Just try working one step at a time. Lets review the equation without the units and just the numbers and variables. \[x*4=(x+4)*3\] Now simplify the right side using the Distributive Property: \[x*4=3x+12\] Now get your like terms to one side: \[4x-3x=12\] Simplify: \[x=12\] Now since your x is just the speed from the beginning of Rider A (the slower rider) You just solve for Rider B's speed since you know Rider A's speed (12mph). Rider B was 4 mph faster than Rider A right? So its: \[x=12\] \[x+4mph\] \[12+4mph=16mph\] The faster rider goes 16mph while the slower rider goes 12mph. :D

OpenStudy (anonymous):

Thanks so much, that really helped!

OpenStudy (anonymous):

Glad to help!